Is (-4,7) a solution to the equation 

3(x - 10) = 6y
Jamie

Answer by AnlytcPhil(1806)

Is (-4,7) a solution to the equation 

3(x - 10) = 6y 

The 1st number of (-4,7) is -4.  So -4 is the number you substitute for x.
The 2nd number of (-4,7) is 7.  So 7 is the number you substitute for y.

In

3(x - 10) = 6y

replace the x by -4:

3(-4 - 10) = 6y

replace the y by (7):

3(-4 - 10) = 6(7)

Combine the numbers inside the parentheses:

3(-14) = 6(7)

Multiply the 3 by the -14

-42 = 6(7)

Multiply the 6 by the 7

-42 = 42

No they aren't equal.  Their absolute values are equal, 
but their signs aren't.  So this is NOT a solution. Next
I'll give you an ordered pair which IS a solution:

--------------------------

Suppose you were given (2,-4) instead.

The 1st number of (2,-4) is 2.  So 2 is the number you substitute for x.
The 2nd number of (2,-4) is -4.  So -4 is the number you substitute for y.

In

3(x - 10) = 6y

replace the x by 2:

3(2 - 10) = 6y

replace the y by (-4):

3(2 - 10) = 6(-4)

Combine the numbers inside the parentheses:

3(-8) = 6(-4)

Multiply the 3 by the -8

-24 = 6(-4)

Multiply the 6 by the (-4)

-24 = -24

These are equal.  So this IS a solution.

Edwin
AnlytcPhil@aol.com